1. A saturated solution of lead(II) chloride was prepared by dissolving PbCl2 solid in water. The concentration of Pb+2 ion in the solution was found to be 1.62*10^-2M . Calculate Ksp for PbCl2 .

2. The value of Ksp for silver chromate, Ag2CrO4 is 9.0*10^-12 . Calculate the solubility of Ag2CrO4 in grams per liter.
Answer #1
Concepts and reason

The concepts used here are based on the solubility product constant and molar solubility.

First, the solubility product of lead chloride is calculated. After that, using the solubility product constant of substance the molar solubility or concentration of silver chromate is calculated.

Fundamentals

Solubility product:

The solubility product helps to determine the solubility of the compound in water.

Higher the value of solubility product, higher will be the solubility of that compound in water. It tells the level of solute dissolves in a particular solution.

Example-

AB,
aA +bB

The expression for the solubility product will be shown below.

K. =[A] [B]
…… (1)

Here, is the concentration of andis the Concentration of , a and b are the coefficients of the respective ion.

Molar solubility:

The solubility (molar solubility) of a solid is expressed as the concentration of the dissolved solid in a saturated solution.

The molar solubility can be calculated as shown below.

aA (s)
bB(aq)+cC(aq)
0 0
K
=xxx

x= K
…… (2)

Here, is the molar solubility.

(1)

The balanced reaction of dissolution of lead chloride is as follows.

PÉCI, (s)
Pb²+ (aq) + 2CF (aq)

The ICE table of dissociation of leadchloride is shown below.

PbCl2 (s)
Pb2+ (aq) + 2C1 (aq)
0
Intial concentration
change in concentration
Equilbrium concentration
-x
+2x
2x

The solubility product expression can be written as follows.

<(+)(x)= dy
{-17[1zqa] = y

To calculate the of РЬСІ,
substitute the value of as 1.62x10 PM
in the above equation.

K. =(1.62x10-²M)(2x1.62x10-?M)
K.=1.62x10-2x(3.24x10-2)
Ko = 1.62x10-2x10.49x10-4
K =1.7x10-5

(2)

The reaction of silver chromate in water is shown below.

Ag, Cro (s)
2Ag* (aq) + Cro - (aq)

The ICE table of dissociation of silver chromate is shown below.

Ag,Cro, (s)
2Ag* (aq) + Cro,?- (aq)
0
Intial concentration
change in concentration
Equilbrium concentration
+2x
2x

The expression for solubility product for silver chromate will be shown below.

[+*o-ɔ][.8v]= *y

Substitute 9.0x10-12
as the value of and as the value of [agt
and as the value of [Cro ²-1
.

9.0x10-12 = (2x)” (x)
9.0x10-12
4
2.25x10-12 = x
1.310x104 mol L=x

Conversion of the unit of solubility from mol per liter to gram per liter is as shown below.

x (g.L-)=Mxx (mol.L-)
…… (3)

Substitute 1.310x104 mol 
as the value of x(mol.L)
and 331.73 g mol
as the value of the molar mass of silver nitrate (M) in equation (3)

x = 331.73 g.mot x1.310x104 mol -
x = 4.34x102 g

Ans: Part 1

The solubility product of lead chloride (PbCI,)
is 1.7x10-5
.